### v2

# CoRanking: Collaborative Ranking with Small and Large Ranking Agents

Abstract
Large Language Models (LLMs) have demonstrated superior listwise ranking performance. However, their superior performance often relies on large-scale parameters (e.g., GPT-4) and a repetitive sliding window process, which introduces significant efficiency challenges. In this paper, we propose CoRanking, a novel collaborative ranking framework that combines small and large ranking models for efficient and effective ranking. CoRanking first employs a small-size reranker to pre-rank all the candidate passages, bringing relevant ones to the top part of the list (e.g., top-20). Then, the LLM listwise reranker is applied to only rerank these top-ranked passages instead of the whole list, substantially enhancing overall ranking efficiency. Although more efficient, previous studies have revealed that the LLM listwise reranker have significant positional biases on the order of input passages. Directly feed the top-ranked passages from small reranker may result in the sub-optimal performance of LLM listwise reranker. To alleviate this problem, we introduce a passage order adjuster trained via reinforcement learning, which reorders the top passages from the small reranker to align with the LLM’s preferences of passage order. Extensive experiments on three IR benchmarks demonstrate that CoRanking significantly improves efficiency (reducing ranking latency by about 70%) while achieving even better effectiveness compared to using only the LLM listwise reranker.
[…]
CoRanking: Collaborative Ranking with Small and Large Ranking Agents
Wenhan Liu1, Xinyu Ma2, Yutao Zhu1, Lixin Su2, Shuaiqiang Wang2 Dawei Yin2  and Zhicheng Dou1thanks: Corresponding author. 1Gaoling School of Artificial Intelligence, Renmin University of China 2Baidu Inc., Beijing, China lwh@ruc.edu.cn, xinyuma2016@gmail.com, dou@ruc.edu.cn
Figure 1: (A) Apply LLM listwise reranker to rerank 100 passages based on sliding window strategy. (B) Utilize a small reranker to pre-rank 100 passages and then apply LLM reranker to rerank top-20 passages. (C) After the reranking process of small reranker, we first use a passage order adjuster to adjust the top-20 passages and then apply the LLM reranker to rerank top-20 passages.
1 Introduction
In recent years, many studies Liu et al. (2024c); Sun et al. (2023); Zhu et al. (2023) have demonstrated strong zero-shot ranking capabilities of large language models (LLMs). A typical approach is listwise ranking which feeds the query and a list of passages into LLM and instructs it to output a reranked list of passage IDs in descending order of relevance. This approach has been widely adopted by many studies, outperforming many ranking approaches Qin et al. (2023); Liang et al. (2022); Sachan et al. (2022) and establishing state-of-the-art performance on major IR benchmarks including TREC Craswell et al. (2020b) and BEIR Thakur et al. (2021) through implementations like RankGPT Sun et al. (2023) using GPT-4.
Despite its effectiveness, listwise ranking faces critical efficiency challenges (Figure 1A), primarily stemming from two factors. First, its state-of-the-art performance relies on massive LLMs (e.g., Qwen2.5-72B Yang et al. (2024) and GPT-4), which introduces significant computational overhead and results in the efficiency issue. Second, to achieve a trade-off between effectiveness and efficiency, listwise ranking usually utilize the sliding window strategy. With fixed window size and step size, this strategy bubbles the relevant passages that are initially lowly ranked to the top. However, the interdependence between windows hinders the parallelization of sliding windows, which results in many repetitive and sequential ranking of sliding window (e.g., totally 9 sliding windows if window size and step size are set as 20 and 10 when reranking 100 passages) and creates an efficiency bottleneck (see Figure 1 (A)).
The sliding window mechanism of LLM reranker primarily serves to elevate initially low-ranked relevant passages into the final reranking window (e.g., top-20). We believe that this repetitive and inefficient process could be replaced by a smaller reranker: By utilizing a smaller reranker to pre-rank the list of passages, the listwise LLM reranker only needs to rerank the passages in the final window. This hybrid approach could dramatically reduce computational costs and enhance the ranking efficiency.
However, existing small rerankers are primarily fine-tuned from relevance labels (human labels Nogueira and Cho (2019) or teacher rankings Pradeep et al. (2023b); Liu et al. (2024c)). We argue that directly feeding the top-ranked passages of these reranker into an LLM listwise reranker may lead to sub-optimal performance (see Figure 1 (B)). This is because the order of input passages significantly impacts the effectiveness of the listwise LLM reranker Sun et al. (2023); Tang et al. (2024), which stems from LLMs’ pre-training biases toward the positions of input tokens. It indicates that the performance of the LLM listwise reranker has a complex relationship with the order of the input passages, rather than solely depending on the relevance ranking of the passages. Thus, the passage order given by small reranker can hardly aligns with the order preferences of LLMs. To address this issue, we introduce a passage order adjuster to rerank the top-ranked passages from the small reranker so that they align with the order preference of the LLM listwise reranker (see Figure 1 (C)).
In this paper, we propose a collaborative ranking framework CoRanking that integrates small and large ranking agents to perform passage ranking in a multi-stage manner. In the first stage, we train a small listwise reranker (SLR) to pre-rank all the passages to elevate the relevant passages to top part of list (e.g., top-20). Different from existing methods Pradeep et al. (2023a, b) that solely use teacher-generated ranking lists as training labels, we propose a human-label-enhanced method to improve the ranking quality. In the second stage, we train a passage order adjuster (POA) using the reinforcement learning algorithm DPO. The POA reorders the top-ranked passages to align with LLM’s order preference. In addition, we design a significance-aware selection strategy $\text{S}^{3}$ to build high-quality preference pairs for DPO training. In the third stage, an LLM listwise reranker is used to further rerank the top-ranked passages, enhancing the overall ranking effectiveness. We conduct extensive experiments on diverse IR benchmarks and the results indicate that our CoRanking framework achieves 70% faster ranking time and even better ranking effectiveness than pure LLM listwise reranker.
Our contributions are summarized as follows:
(1) We propose a ranking framework CoRanking that enables small and large ranking agents to collaborate with each other, thereby achieving efficient and effective passage ranking.
(2) We introduce a novel passage order adjuster (POA) that aligns the the order of reranked passages from small reranker with the preference of large listwise reranker.
(3) We design the $\text{S}^{3}$ strategy which build valuable training samples for the DPO algorithm during POA’s training, improving the effectiveness of preference alignment.
Figure 2: Our collaborative ranking framework CoRanking consists of three stages. In the first stage, SLR reranks top-100 passages based on sliding window strategy. In the second stage, POA reranks the top-20 passages to align with the positional preference of LLR. In the final stage, LLR is utilized to rerank the top-20 passage to obtain the final ranking list. The training details of SLR and POA are shown in yellow and cyan parts respectively.
2 Methodology
In this section, we present a collaborative ranking framework CoRanking that integrates both small and large passage rerankers. As illustrated in Figure 2, the framework comprises three key components: (1) a small listwise reranker (SLR) for preliminary passage reranking, (2) a passage order adjuster (POA) that aligns the order of SLR’s top-ranked passages with the LLM listwise reranker’s preference, and (3) an LLM listwise reranker (LLR) for final ranking. We begin by formalizing the passage ranking task, then elaborate on each component’s design.
2.1 Task Definition
The objective of passage ranking is to rerank a list of retrieved passages $[p_{1},\ldots,p_{N}]$ based on their relevance to a query $q$. The listwise ranking approach takes both $q$ and a set of passages as input, and generates a reranked sequence of passage IDs (e.g., [4][2]…). Due to LLMs’ input length constraints, existing listwise ranking approaches typically employ a sliding window strategy that iteratively processes passage subsets. Specifically, this strategy applies a window of size $w$ that moves from the end to the beginning of the passage list with a step size $s$, dynamically promoting relevant passages to higher positions. Following previous studies Sun et al. (2023); Liu et al. (2024c), we set the total number of passages $N$ to 100, with window size $w$ and step size $s$ set to 20 and 10, respectively.
2.2 Stage1: Small Listwise Reranker (SLR)
To mitigate the latency and computational overhead of the LLM listwise reranker (LLR), we propose using a small listwise reranker (SLR) to pre-rank the entire passage list. The SLR aims to elevate relevant passages into the final sliding window (i.e., the top-20 passages), thereby reducing the number of LLR’s sliding window iterations. Here we adopt listwise training because we find that it performs better than pointwise training and the SFT-trained SLR naturally serves as the initialization for our listwise passage order adjuster, which will be detailed in Section 2.3.
Training Label. Conventional listwise reranker training typically employs knowledge distillation Pradeep et al. (2023a, b), which uses a strong teacher model (e.g., GPT-4) to rerank BM25-retrieved top-20 passages, with the teacher-generated rankings serving as supervised fine-tuning (SFT) labels. However, the teacher model can not perform consistently well across all queries (e.g., long-tail queries), which can result in suboptimal performance on these queries. To further improve the SFT label, we propose a human-label-enhanced ranking construction (HRC) method that improves teacher-generated rankings with human-label.
Specifically, we first iteratively rerank BM25-retrieved top-20 passages $M$ times using the teacher model in a listwise manner, where the $i$-th reranking input is the output ranking from the $(i-1)$-th iteration. This progressive reranking approach, motivated by prior studies Pradeep et al. (2023b, a), enables gradual improvement of ranking lists. We select the iteration where human-annotated relevant passages are ranked at the top (NDCG=1) as the gold ranking.
Training Loss. After obtaining the gold ranking, we concatenate the query $q$ and the candidate passages $[p_{1},\ldots,p_{N}]$ together as the training input $x$ and optimize our SLR by minimizing the standard language modeling loss $\mathcal{L}$:
$\mathcal{L}$ $=-\sum_{i=1}^{|y|}\log(P_{\theta}(y_{i}\mid x,y_{<i})),$ (1)
$x$ $=q\circ\text{[1]}\circ p_{1}\circ\text{[2]}\circ p_{2}\ldots\text{[N]}\circ p_{N}$ (2)
where $[i]$ represents the ID of passage $p_{i}$ and $y$ is the gold ranking label (e.g., “[4] [2] …”).
2.3 Stage2: Passage Order Adjuster (POA)
While our SLR is trained to rank relevant passages at the top of the list, the order of top-ranked passages (i.e., the top-20 passages) may not align well with the positional preferences of LLR. To address this mismatch, we employ reinforcement learning (RL) algorithm to train a passage order adjuster POA based on LLR’s feedback to rerank these passages. We adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm for its optimization stability. The following sections elaborate on the construction of DPO training data and model optimization.
Training Data. The training data of DPO consists of a series of prediction pairs. Each prediction pair contains a winner and a loser prediction, namely passage ID lists. The reward $R$ for each prediction $L$ is computed by feeding the corresponding passage list to LLR and evaluating the ranking metric $M$ of LLR’s output ranking $L^{\prime}$ based on human-annotated label $y$:
$R=M(L^{\prime},y),$ (3)
where we use NDCG@10 as our metric $M$.
We apply our trained SLR to rerank BM25-retrieved top-100 passages for training queries and use the top-20 reranked passages as training candidate passages, which aims to align with the input passage distribution encountered by POA during inference. To generate the preference pairs for training POA, we first sample $m$ predictions from SLR by taking the candidates as the input, then evaluate these predictions through Equation 3. After that, the highest-reward list is selected as the winner prediction, while others serve as loser predictions. Next, we pair the winner prediction with all the loser predictions to construct initial training preference pairs.
To ensure the quality of these preference pairs, we introduce a significance-aware selection strategy $\text{S}^{3}$ to filter out low-quality pairs. Specifically, the $\text{S}^{3}$ strategy follows two rules: (1) Significance of Winner: The reward of the winner prediction must be 1 and exceed LLR’s ranking metric (i.e., the NDCG@10 of directly use LLR to rerank the candidate passages), ensuring the winner prediction provides positive improvement for final reranking of LLR; (2) Significant Gap between winner and loser: The reward difference between winner prediction and loser prediction must exceed a threshold $\mu$, eliminating noisy pairs with marginal differences.
Optimization. Given constructed preference pairs, we optimize POA using the following DPO objective:
$\mathcal{L}_{\text{DPO}}$ $=-\underset{x,y_{w},y_{l}}{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}^{w}\pi_{f}^{t}}{\pi_{f}^{w}\pi_{\theta}^{t}}\right)\right],$ (4)
$\pi_{\theta}^{w}$ $=\pi_{\theta}(y_{w}\mid x),\quad\pi_{f}^{t}=\pi_{f}(y_{l}\mid x),$ (5)
$\pi_{f}^{w}$ $=\pi_{f}(y_{w}\mid x),\quad\pi_{\theta}^{t}=\pi_{\theta}(y_{l}\mid x),$ (6)
where $x$ represents the training input whose format is the same as Equation 1 and $y_{w}$, $y_{l}$ represent the winner prediction and loser prediction, respectively. $\pi_{\theta}$ is the policy model to be optimized and $\pi_{f}$ is the original policy model which is initialized with our trained SLR. The frozen $\pi_{f}$ acts as a regularizer to maintain generation stability during optimization and $\beta$ balances preference alignment and regularization towards $\pi_{f}$.
2.4 Stage3: LLM Listwise Reranker (LLR)
After the Passage Order Adjuster (POA) adjusts the top-20 passages, we use the LLM Listwise Reranker (LLR) to rerank these 20 passages. This avoids LLR’s inefficient sliding window operations on all 100 passages, significantly improving ranking efficiency. The specific prompt template for LLR is provided in Appendix A, which is consistent with previous studies Pradeep et al. (2023b); Liu et al. (2024c).
3 Experiments
3.1 Setting
Evaluation Datasets
We evaluate our framework on three established information retrieval benchmarks: TREC DL Craswell et al. (2020b), BEIR Thakur et al. (2021), and BRIGHT Su et al. (2024). For the TREC DL benchmark, we adopt the TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets derived from MS MARCO V1. The BEIR benchmark Thakur et al. (2021) is selected to evaluate cross-domain generalization capabilities. We choose three representative datasets: TREC-Covid, Robust04, and Trec-News. BRIGHT Su et al. (2024) is a reasoning-intensive IR benchmark whose relevance assessment transcends simplistic keyword matching or semantic similarity, instead demanding intentional and deliberate reasoning. We choose Economics, Earth Science and Robotics datasets for evaluation. The detailed descriptions of each dataset will be provided in Appendix B.
TREC DL BEIR BRIGHT
Models Strategy DL19 DL20 Covid Robust04 News Econ. Earth. Rob. Avg.
BM25 - 50.58 47.96 59.47 40.70 39.52 16.45 27.91 10.91 36.69
Single Reranker
monoBERT (340M) Pointwise 70.46 67.55 70.85 46.11 47.61 10.45 19.98 11.34 43.04
monoT5 (3B) Pointwise 71.57 69.83 72.39 49.81 49.34 9.65 19.16 8.56 43.79
RankT5 (3B) Pointwise 71.83 70.27 80.19 51.45 49.11 8.66 24.75 13.95 46.28
Qwen-pointwise (3B) Pointwise 70.62 64.99 81.77 53.60 47.78 14.13 26.25 9.89 46.13
RankVicuna (7B) Sliding 67.72 65.98 79.19 48.33 47.15 11.05 16.30 8.03 42.97
[…]
SLR (3B) Sliding 70.24 68.85 81.85 53.68 49.85 15.56 29.47 13.68 47.90
[…]
Collaborative Rerankers
[…]
Table 1: Results (NDCG@10) on TREC, BEIR and BRIGHT benchmarks. The “Pointwise”, “Sliding”, ’Full’ and “Collaborative” in “Strategy” column represent pointwise ranking, sliding window strategy, full ranking Liu et al. (2024c) and collaborative ranking, respectively. The best and second-best results for “Single Reranker” part and “Collaborative Rerankers” part are indicated in bold and underlined, respectively. “Avg.” represents the averaged result of all 8 datasets.
Implementation Details
As for the supervised fine-tuning of our SLR, we use Qwen2.5-3B-Instruct as the backbone model. The training queries consists of totally 2k queries sampled from MS MARCO training set. We utilize Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4 as our teacher model to SFT our SLR and also employ it as the LLR model throughout this study unless otherwise specified. Unlike previous studies Pradeep et al. (2023b); Liu et al. (2024c) that adopted GPT-series models (e.g., GPT-4o) as the teacher model, we select Qwen2.5-72B-Instruct as it demonstrates comparable and even better performance to GPT-4o across most datasets while being more cost-effective. Nevertheless, we have included additional experimental results using GPT-4o as the LLR model in Section 3.6. We train SLR for 4 epochs with a learning rate of 5e-6 and batch size of 1. As for the RL training of POA, we use our trained SLR for model initialization. We set the threshold $\mu$ as 0.4 and generate about 12k preference pairs based on 2k queries for DPO algorithm. The batch size, learning rate and the hyper-parameter $\beta$ of DPO are set as 1, 1e-6 and 0.4 respectively.
Baselines
We compare our framework against a series of baselines, which can be divided into two parts: (1) Single reranker, (2) Collaborative rerankers (i.e.reranking with small and large rerankers).
As for single reranker, we compare our trained SLR with a series of state-of-the-art supervised rerankers, including monoBERT (340M) Nogueira and Cho (2019), monoT5 (3B) Nogueira et al. (2020), RankT5 (3B) Zhuang et al. (2023b), RankVicuna (7B) Pradeep et al. (2023a), RankZephyr (7B) Pradeep et al. (2023b) and $\textbf{RankMistral}_{\text{100}}$. The first three are pointwise rerankers trained from human labels in MS MARCO dataset. The last three are listwise rerankers which are distilled from GPT-3.5, GPT-4 and GPT-4o, respectively. To ensure more fair comparison with SLR in model architecture and training data, we introduce baseline Qwen-pointwise which is a pointwise reranker based on Qwen2.5-3B-Instruct. This model concatenates query-passage pair as input and outputs a relevance score. We utilize the same gold ranking label as SLR and optimize the model with RankNet loss, which is the same as what  Sun et al. (2023) did. We also report the ranking performance of our SLR and LLR which reranks all the passages using sliding window strategy.
As for collaborative rerankers, we introduce three baselines, including monoT5 (3B) + LLR, RankT5 (3B) + LLR and Qwen-pointwise (3B) + LLR. The three baselines first use corresponding small reranker (3B) to rerank all the passages and then apply LLR to rerank top-20 passages.
3.2 Overall Results
The main experimental results are presented in Table 1, from which we draw the following observations:
(1) Our collaborative reranking framework CoRanking achieves the best average performance (51.39) across all datasets, outperforming both single rerankers and other collaborative baselines. Notably, it even surpasses the LLR by about 0.5 points on average. These results demonstrate that our framework could effectively align the passage order generated by the SLR (3B) with the preferences of the large language model (LLR), thereby achieving better ranking performance than only using LLM reranker LLR.
(2) Regarding single rerankers, our SLR (3B) achieves an average score of 47.90, outperforming all the pointwise rerankers including monoBERT, monoT5, RankT5 and Qwen-pointwise, and even surpassing 7B-scale RankVicuna and RankZephyr. Furthermore, SLR surpasses Qwen-pointwise by about 1.8 average points under identical training conditions, proving listwise reranking’s superiority through comparison of multiple passages for precise relevance assessment.
(3) Collaborative reranker baselines (e.g., RankT5 (3B) + LLR) improve over corresponding single rerankers (e.g.RankT5 (3B)). However, these baselines still lags behind and LLR and our CoRanking, indicating that directly feeding small rerankers’ ranking results to LLR fails to align with LLR’s order preference.
Model TREC Avg. BEIR Avg. BRIGHT Avg.
CoRanking 71.64 64.96 24.32
w/o $\text{S}^{3}$ strategy 70.10 63.86 23.26
w/o POA 70.23 63.80 22.74
SLR 69.55 61.79 19.57
w/o HRC 68.22 59.58 19.43
Table 2: Average performance (NDCG@10) of ablated models on TREC (DL19, DL20), BEIR (Covid, Robust04, News), and BRIGHT (Economics, Earth Science, Robotics).
3.3 Ablation Study
To evaluate the contribution of each component in our framework, we conduct comprehensive ablation experiments. Table 2 presents the average performance for each individual benchmark, with detailed per-dataset results available in Table 4.
First, to evaluate our proposed $\text{S}^{3}$ sampling strategy for generating the preference pairs of DPO, we introduce a variant, denoted as “w/o $\text{S}^{3}$”. This variant uses random sampling for preference pair selection, without ensuring the significant reward of winner prediction and the reward gap between the winner and loser prediction. The performance degradation across all benchmarks (e.g., 1.54-point drop on TREC Avg.) demonstrate the effectiveness of our $\text{S}^{3}$ sampling in identifying valuable preference pairs for effective DPO training.
We further remove our POA to validate its alignment effectiveness, denoted as “w/o POA”. In this case, the LLR directly reranks the top-20 passages after SLR’s reranking. The notable performance decline (especially 1.6-point decrease on BRIGHT Avg.) demonstrates that POA effectively align with the passage order preference of LLR.
Finally, to validate our human-label-enhanced ranking construction (HRC) for fine-tuning SLR, we create a variant (w/o HRC) which generates ranking label by utilizing LLR to rerank training passages only one time and choosing not to use human label for filtering high-quality ranked lists. The significant performance drops across benchmarks (especially BEIR) highlight HRC’s effectiveness on generating high-quality ranking labels.
3.4 Efficiency Analysis
As mentioned in Section 1, our primary target of CoRanking is to mitigate the efficiency issue of pure LLM-based reranking. To quantitatively demonstrate this advantage, we compare our CoRanking against LLR across all 8 datasets. Our experiments are conducted on 8×40GB NVIDIA A100 GPUs using the vLLM inference framework [^2] https://github.com/vllm-project/vllm. We average across all queries within each dataset and show the comparison results in Figure 3.
Across all datasets, our CoRanking approach demonstrates a significant reduction in latency, achieving a 67%-74% decrease compared to the LLR. This substantial improvement highlights the efficiency advantage of CoRanking. Additionally, the latency of less than 1 second achieved by CoRanking across most datasets also showcases its practical efficiency in real-world search engines. Lastly, the ranking latency varies significantly across different datasets, which is primarily due to the differences in passage lengths. This is because longer passages inherently require more computational time for processing.
Figure 3: Ranking latency (seconds/query) of LLR and CoRanking on 8 datasets.
Figure 4: The changes of NDCG@10 after reranking by SLR, POA, and LLR.
3.5 Detailed Analysis of POA
Our POA is designed to adjust the order of top-20 passages reranked by SLR to align them more with the preferences of the LLR. However, whether this adjustment improves or degrades the ranking quality of top-20 passages remains unknown. We believe this is an interesting question because it helps us understand the reasons why POA works.
In this part, we selected four datasets—DL19, Robust04, News, and Economics—and reported the changes in the ranking metric NDCG@10 after ranking by SLR, POA, and LLR respectively. We also report the performance of the ablated variant SLR+LLR (i.e., directly using LLR to rerank top-20 passages from SLR without POA’s adjustment) as a baseline. The results are shown in Figure 4. After POA’s adjustment, the ranking metric on three datasets—DL19, Robust04, and News—decreases (except for Economics dataset). This indicates that a passage ranking list with high-ranking metric does not necessarily align with LLR’s preferences. This may be due to the positional bias of input tokens during the pre-training of LLMs, as revealed in previous studies Tang et al. (2024); Liu et al. (2024a).
Model DL19 News Earth. Avg.
BM25 50.58 39.52 27.91 39.34
Using GPT-4o as LLR
RankT5 (3B) + LLR 73.17 49.31 29.64 50.71
Qwen-pointwise (3B) + LLR 72.35 48.36 34.62 51.78
SLR+LLR 73.10 48.75 34.97 52.27
CoRanking 74.10 50.19 35.93 53.41
Using DeepSeek-V3 as LLR
RankT5 (3B) + LLR 71.90 49.47 31.47 50.95
Qwen-pointwise (3B) + LLR 72.16 49.07 35.38 52.20
SLR+LLR 72.62 50.91 38.09 53.87
CoRanking 73.79 51.48 39.01 54.76
Table 3: The performance (NDCG@10) of CoRanking based on different LLRs. “Avg.” represents the average performance across the three datasets.
3.6 Generalization on Other LLMs
In the previous experiments, we use Qwen2.5-72B-Instruct as our LLR model. However, it remains unclear whether our proposed CoRanking framework can generalize effectively to other LLR models. To evaluate this, we conducted additional experiments with two different LLRs: GPT-4o [^3] The version of GPT-4o we used is gpt-4o-2024-08-06. and DeepSeek-V3, and compare our CoRanking with three collaborative rerankers (“RankT5 (3B) + LLR”, “Qwen-pointwise (3B) + LLR”, and “SLR + LLR”). Note that “SLR + LLR” represents using LLR to directly rerank the top-20 passages reranked by SLR without POA. For evaluation, we select one dataset from each of the three benchmarks: DL19 from TREC, News from BEIR, and Earth Science from BRIGHT. As illustrated in Table 3, our CoRanking consistently outperforms the best baseline “SLR + LLR” by about 1 point on average, regardless of the LLR used. This demonstrates the positional preference alignment of our POA could generalize to different LLMs.
4 Related Work
4.1 LLMs for Passage Ranking
The application of large language models (LLMs) into retrieval systems Zhu et al. (2023) has driven substantial research efforts in passage ranking tasks. Current LLM-based ranking methods can be categorized into three categories: pointwise, pairwise, and listwise. Pointwise methods Liang et al. (2022); Sachan et al. (2022); Zhuang et al. (2023a); Liu et al. (2024e, d) compute the relevance score for each query-passage pair. Pairwise methods Qin et al. (2023); Luo et al. (2024) compare two passages at a time and identify the more relevant one. Listwise methods Sun et al. (2023); Pradeep et al. (2023a); Reddy et al. (2024); Liu et al. (2024c, b); Chen et al. (2024) directly take a list of passages as the input and output the passage identifiers in descending order of their relevance. Through comparing multiple passages, listwise reranker could assess the relevance more comprehensively and yield better ranking results. Due to the superior performance of listwise ranking Sun et al. (2023), it has attracted the interest of many researchers. For example,  Tang et al. (2024) propose a permutation self-consistency method to address the sensitivity of listwise ranking to initial passage order and achieve a more accurate and positionally unbiased ranking. Besides,  Pradeep et al. (2023a) and  Pradeep et al. (2023b) propose to distill the knowledge of strong listwise reranker (e.g., GPT-4) into moderate-sized models to enhance ranking efficiency while reducing computational overhead.
The superior performance of LLM listwise rerankers is often attributed to the utilization of large-scale LLMs and the sliding window strategy, which leads to efficiency issue in ranking. In this paper, we introduce a novel framework that leverages the collaborative ranking of large-size rerankers and small-size rerankers to enhance the ranking efficiency while maintaining ranking effectiveness.
4.2 Preference Alignment
Preference alignment has emerged as a critical research direction, aiming to align model behaviors with human or task-specific preferences. Early approaches primarily relied on reinforcement learning from human feedback (RLHF), where models are fine-tuned using reward signals derived from human preference annotations. To reduce the computational complexity of RLHF, recent studies proposed more efficient alternatives. Direct Preference Optimization (DPO) Rafailov et al. (2023) reformulates preference learning as a supervised objective, bypassing explicit reward modeling while achieving competitive performance. Recent extensions like Step-DPO Lai et al. (2024) further decompose preference alignment into multi-step optimization. Besides, Dong et al. (2023) introduced rejection sampling fine-tuning to align LLMs with human preferences without requiring reinforcement learning pipelines. In this paper, to address the positional sensitivity of LLM listwise reranker, we propose training a passage order adjuster with the DPO algorithm to align the passage order generated by the small reranker with the positional preferences of the LLM-based listwise reranker.
5 Conclusion
In this paper, we introduce a novel collaborative ranking framework CoRanking which combines small and large rerankers for more efficient and effective passage ranking. Specifically, we first train a small listwise reranker SLR to move the relevant passages to the top of the list. Then we introduce a passage order adjuster POA to reorder the top-ranked passages to better align with the positional preference of LLM listwise reranker. Finally, the LLM listwise reranker is utilized to solely rerank the top passages. Extensive experiments demonstrate the efficiency advantage and superior performance of CoRanking.
Limitations
We acknowledge some limitations in our work. Firstly, our framework consists of three models, making its online deployment more complex compared to using a single unified reranker. Secondly, we only use Qwen2.5-3B-Instruct as the backbone model of our SLR and POA. We believe it is meaningful to attempt using different backbones to demonstrate the effectiveness of our CoRanking framework.
Appendix A Listwise Ranking Prompt
The listwise ranking prompt of LLR used in this paper is shown as bellow:
Prompt: Listwise Ranking Prompt You are RankLLM, an intelligent assistant that can rank passages based on their relevancy to the query. I will provide you with {num} passages, each indicated by a numerical identifier []. Rank the passages based on their relevance to the search query: {query}. [1] {passage 1} [2] {passage 2} [{num}] {passage {num}} Search Query: {query}. Rank the {num} passages above based on their relevance to the search query. All the passages should be included and listed using identifiers, in descending order of relevance. The output format should be [] > [], e.g., [4] > [2]. Only respond with the ranking results, do not say any word or explain.
Appendix B Datasets Details
In this part, we provide the details of each evaluation dataset we used:
$\bullet$ DL19/DL20 The TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets are derived from the MS MARCO V1 passage ranking corpus. DL19 contains 43 test queries with dense relevance annotations from the TREC 2019 Deep Learning Track, while DL20 includes 54 queries from the 2020 edition. Both datasets focus on web search scenarios and are widely adopted as standard benchmarks for large-scale information retrieval systems.
$\bullet$ TREC-Covid Voorhees et al. (2020): A biomedical dataset comprising scientific articles related to COVID-19, with 50 test queries from the TREC 2020 COVID Track.
[…]
Appendix C Baselines Details
In this paper, we include two kinds of baselines: (1) Single reranker and (2) Collaborative rerankers. The details of single reranker are shown as below: $\bullet$ monoBERT (340M) Nogueira and Cho (2019): A BERT-large-based cross-encoder re-ranker.
$\bullet$ monoT5 (3B) Nogueira et al. (2020): A T5-based sequence-to-sequence re-ranker that computes relevance scores via pointwise ranking loss.
$\bullet$ RankT5 (3B) Zhuang et al. (2023b): A T5-based re-ranker optimized with listwise ranking loss.
$\bullet$ Qwen-pointwise (3B): To ensure fair comparison in model architecture and training data, we fine-tune a pointwise re-ranker based on Qwen2.5-3B-Instruct. This model concatenates queries and passages as input to predict relevance scores, trained on SLR data with RankNet loss following Sun et al. (2023).
$\bullet$ RankVicuna Pradeep et al. (2023a): RankVicuna is a listwise re-ranker distilled from GPT-3.5-generated ranked lists.
$\bullet$ RankZephyr Pradeep et al. (2023b): RankZephyr is a listwise reranker distilled from GPT-3.5 and GPT-4.
$\bullet$ LLR (72B): Our teacher model that employs a zero-shot sliding window strategy for passage re-ranking.
TREC BEIR BRIGHT
Model DL19 DL20 Covid Robust04 News Econ. Earth. Rob. Avg.
CoRanking 72.79 70.48 84.90 57.71 52.26 22.72 31.20 19.03 51.39
       w/o $\text{S}^{3}$ strategy 70.90 69.30 83.18 56.91 51.48 21.88 30.38 17.51 50.19
       w/o POA 70.34 70.11 84.14 56.99 50.26 21.20 30.50 16.51 50.01
SLR 70.24 68.85 81.85 53.68 49.85 15.56 29.47 13.68 47.90
       w/o HRC 69.53 66.90 79.36 50.61 48.77 15.59 29.66 13.04 46.68
Table 4: Results (NDCG@10) of ablated models on TREC, BEIR, and BRIGHT benchmarks. The best-performing models are marked in bold.