### v3

# Wenhan Liu1, Xinyu Ma2, Yutao Zhu1, Lixin Su2, Shuaiqiang Wang2 Dawei Yin2 Zhicheng Dou1

Abstract
Listwise ranking based on Large Language Models (LLMs) has achieved state-of-the-art performance in Information Retrieval (IR). However, their effectiveness often depends on LLMs with massive parameter scales and computationally expensive sliding window processing, leading to substantial efficiency bottlenecks. In this paper, we propose a Collaborative Ranking framework (CoRanking) for LLM-based listwise ranking. Specifically, we strategically combine an efficient small reranker and an effective large reranker for collaborative ranking. The small reranker performs initial passage ranking, effectively filtering the passage set to a condensed top-k list (e.g., top-20 passages), and the large reranker (with stronger ranking capability) then reranks only this condensed subset rather than the full list, significantly improving efficiency. We further address that directly passing the top-ranked passages from the small reranker to the large reranker is suboptimal because of the LLM’s strong positional bias in processing input sequences. To resolve this issue, we propose a passage order adjuster learned by RL that dynamically reorders the top passages returned by the small reranker to better align with the large LLM’s input preferences. Our extensive experiments across three IR benchmarks demonstrate that CoRanking achieves superior efficiency, reducing ranking latency by approximately 70% while simultaneously improving effectiveness, compared to the standalone large reranker.
[…]
[Uncaptioned image] CoRanking: Collaborative Ranking with Small and Large Ranking Agents
Wenhan Liu1, Xinyu Ma2, Yutao Zhu1, Lixin Su2, Shuaiqiang Wang2 Dawei Yin2  and Zhicheng Dou1thanks: Corresponding author. 1Gaoling School of Artificial Intelligence, Renmin University of China 2Baidu Inc., Beijing, China lwh@ruc.edu.cn, xinyuma2016@gmail.com, dou@ruc.edu.cn
Figure 1: (A) Large Listwise Reranker: Apply a large listwise reranker to rerank 20 passages from 100 candidates based on a sliding window strategy. (B) Naive Collaboration: Pre-rank 100 passages with a small listwise reranker using a sliding window strategy, then rerank the top 20 with a large listwise reranker. (C) CoRanking: Pre-rank 100 passages with a small reranker using a sliding window strategy, adjust the order of the top 20 via the passage order adjuster, then rerank them with the large listwise reranker. “Large/Small Reranker Calls” represents the number of sliding windows for listwise ranking.
1 Introduction
In recent years, many studies Liu et al. (2024c); Sun et al. (2023); Zhu et al. (2023) have demonstrated strong zero-shot passage ranking capabilities of large language models (LLMs). A typical approach is listwise ranking, which feeds the query and a list of passages into an LLM and instructs it to output a ranked list of passage IDs in descending order of relevance. This approach has established state-of-the-art performance Sun et al. (2023) on major IR benchmarks, including TREC Craswell et al. (2020b) and BEIR Thakur et al. (2021).
Despite its effectiveness, LLM listwise passage ranking faces critical efficiency challenges, as illustrated in Figure 1(A), primarily stemming from two factors: First, its state-of-the-art performance relies on LLMs with massive parameters (such as Qwen2.5-72B or GPT-4), which introduces significant computational overhead and results in an efficiency issue. Second, due to the limitation of the context length of LLMs, existing listwise ranking methods typically employ a sliding window strategy Sun et al. (2023). With a fixed window size (e.g., 20 passages) and step size (e.g., 10 passages), this strategy bubbles the relevant passages that are initially lowly ranked to the top. However, the interdependence between windows hinders the parallelization of sliding windows, which results in many repetitive and sequential rankings of sliding windows (e.g., totally 9 sliding windows if window size and step size are set as 20 and 10 when ranking the top 20 from 100 passages) and creates an efficiency bottleneck (especially when employing LLMs with massive parameters).
The sliding window strategy of LLM reranker primarily serves to elevate initially low-ranked relevant passages into the final reranking window (e.g., top 20). We believe that this repetitive and inefficient process could be replaced by a small listwise reranker. [^1] In this work, we consider models with parameters less than or equal to 3B as small rerankers. By utilizing a small listwise reranker to pre-rank the whole list of passages (e.g., 100 passages) using a sliding window strategy, the large listwise reranker, which has stronger ranking capability, only needs to rerank the passages in the final window (i.e., containing the top 20 passages). This hybrid approach could dramatically reduce computational costs and improve efficiency.
However, existing small listwise rerankers are mainly trained on golden lists whose passages are ranked solely based on relevance Pradeep et al. (2023b, a). We argue that directly feeding the top-ranked passages returned by these small listwise rerankers into a large listwise reranker may lead to sub-optimal performance (see Figure 1 (B)). This is because the effectiveness of a listwise reranker has a complex relationship with the order of input passages, rather than being determined exclusively by their relevance, as indicated by existing studies Sun et al. (2023); Tang et al. (2024). For example,  Tang et al. (2024) reveal that relevant passages appearing in the middle of the list tend to be ranked at the bottom by the listwise reranker. This stems from LLMs’ pre-training biases toward the positions of input tokens Liu et al. (2024a). While the passage order generated by the small listwise reranker, which is trained using relevance-based ranking, may deviate from the large listwise reranker’s order preference. To address this issue, we introduce a passage order adjuster (also a small listwise reranker) to rerank the top-ranked passages from the small listwise reranker so that they align with the order preference of the large listwise reranker (see Figure 1 (C)).
More specifically, we propose a collaborative ranking framework, namely CoRanking, that integrates small and large ranking models to perform passage ranking in a multi-stage manner. In the first stage, we train a small listwise reranker (SLR) to pre-rank all the passages to elevate the relevant passages to the top part of the list (e.g., top-20). Different from existing methods Pradeep et al. (2023a, b) that solely use teacher-generated ranking lists as training labels, we propose a human-label-enhanced method to improve the ranking quality. In the second stage, we train a passage order adjuster (POA) using the reinforcement learning algorithm DPO. The POA reorders the top-ranked passages to align with the large listwise reranker’s order preference. In addition, we design a significance-aware selection strategy $\text{S}^{3}$ to build high-quality preference pairs for DPO training. In the third stage, a large listwise reranker is used to further rerank the top-ranked passages, enhancing the overall ranking effectiveness. Extensive experiments on diverse IR benchmarks demonstrate that our CoRanking framework achieves a 70% reduction in ranking latency while maintaining superior effectiveness, compared to a standalone large listwise reranker.
Our contributions are summarized as follows:
(1) We propose a ranking framework, CoRanking, that enables small and large LLMs to collaborate, thereby achieving efficient and effective passage ranking.
(2) We propose a novel passage order adjuster (POA) that aligns the output passages of the small listwise reranker with the order preference of the large listwise reranker. We further design the $\text{S}^{3}$ strategy, which builds valuable preference pairs for POA’s training.
(3) We conduct extensive experiments on diverse IR benchmarks, and the results demonstrate that our CoRanking achieves a significant efficiency improvement compared to a standalone large listwise reranker while exhibiting superior performance.
Figure 2: The “Inference” part shows how our CoRanking works with both large and small listwise rerankers. As for training of the small listwise reranker, we design a human-label-enhanced ranking construction method to generate gold rankings for supervised fine-tuning. As for training the passage order adjuster, we proposed the $\text{S}^{3}$ strategy to obtain high-quality preference pairs for DPO reinforcement learning.
2 Methodology
In this section, we present a collaborative ranking framework CoRanking that integrates both small and large passage rerankers. As illustrated in Figure 2, the framework comprises three key components: (1) a small listwise reranker, (2) a passage order adjuster, and (3) a large listwise reranker. Next, we will introduce the definition of the ranking task, the inference process of CoRanking, and the training details.
2.1 Task Definition
The aim of passage ranking is to rerank a list of retrieved passages $p_{1},\ldots,p_{N}]$ based on their relevance to a query $q$. The listwise ranking approach takes both $q$ and a set of passages as input, and generates a reranked sequence of passage IDs (e.g., [4] [2]…). Due to the constraints of LLMs’ input length, existing listwise ranking approaches typically employ a sliding window strategy that iteratively processes passage subsets. Specifically, this strategy applies a window of size $w$ that moves from the end to the beginning of the passage list with a step size $s$, dynamically promoting relevant passages to higher positions. Following previous studies Sun et al. (2023), we set the total number of passages $N$ to 100, with window size $w$ and step size $s$ set to 20 and 10, respectively.
2.2 Inference Process of CoRanking
Given the passages to rerank, our CoRanking (see Figure 2) first applies the small listwise reranker to pre-rank all the passages using a sliding window strategy. Then, we use the passage order adjuster to rerank the top-20 passages in a listwise manner, so that their order meets the preference of the large listwise reranker. Finally, the large listwise reranker is applied to further rerank the top-20 passages to obtain the final ranking list. This avoids the large listwise reranker’s inefficient sliding window operations on all 100 passages, significantly improving ranking efficiency. The specific prompt template for a large listwise reranker is provided in Appendix A, which is consistent with previous studies Pradeep et al. (2023b); Liu et al. (2024c). In the following, we will illustrate the training details of our small listwise reranker and passage order adjuster.
2.3 Small Listwise Reranker (SLR)
To mitigate the latency and computational overhead of the large listwise reranker (denoted as LLR), we propose using a small listwise reranker (denoted as SLR) to pre-rank the entire passage list. The SLR aims to elevate relevant passages into the final sliding window (i.e., the top-20 passages), thereby reducing the number of LLR’s sliding window iterations. Here we adopt listwise training because we find that it performs better than other approaches (e.g., pointwise training Nogueira and Cho (2019)), and the SFT-trained SLR naturally serves as the initialization for our listwise passage order adjuster, which will be detailed in Section 2.4.
Training Label Conventional listwise reranker training typically employs knowledge distillation Pradeep et al. (2023a, b), which uses a strong teacher model (e.g., GPT-4) to rerank BM25-retrieved top-20 passages, with the teacher-generated rankings serving as supervised fine-tuning (SFT) labels. However, the teacher model can not perform consistently well across all queries, such as long-tail queries, which can result in suboptimal performance. To further improve the SFT label, we propose a human-label-enhanced ranking construction (HRC) method that improves teacher-generated rankings with human labels.
Specifically, we first iteratively rerank BM25-retrieved top-20 passages $M$ times using a powerful teacher model in a listwise manner, where the $(i+1)$-th reranking input is the output ranking from the $i$-th iteration. In this paper, we apply LLR as our teacher model and set the maximum iteration number to 5. This progressive reranking approach is motivated by prior studies Pradeep et al. (2023b, a) which reveal that iteratively reranking enables gradual improvement of ranking lists. Then, we choose the ranking where human-annotated relevant passages are ranked at the top (i.e., NDCG=1) during the iteration process as the gold ranking.
Training Loss After obtaining the gold ranking, we concatenate the query $q$ and the candidate passages $p_{1},\ldots,p_{N}]$ together as the training input $x$ and optimize our SLR by minimizing the standard language modeling loss $\mathcal{L}$:
$\mathcal{L}$ $=-\sum_{i=1}^{|y|}\log(P_{\theta}(y_{i}\mid x,y_{<i})),$ (1)
$x$ $=q\circ\text{[1]}\circ p_{1}\circ\text{[2]}\circ p_{2}\ldots\text{[N]}\circ p_{N},$ (2)
where $[i]$ represents the ID of passage $p_{i}$ and $y$ is the gold ranking label (e.g., “[4] [2] …”).
2.4 Passage Order Adjuster (POA)
While our SLR is trained to rank relevant passages at the top of the list, the order of top-ranked passages (i.e., the top-20 passages) may not align well with the order preference of LLR. To address this issue, we employ a reinforcement learning (RL) algorithm to train a passage order adjuster POA based on LLR’s feedback to rerank these top-ranked passages. We adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm for its optimization stability. The following sections elaborate on the construction of DPO training data and model optimization.
Training Data The training data of DPO consists of a series of prediction pairs. Each prediction pair contains a winner and a loser prediction, namely, a passage ID list. The reward $R$ for each prediction $L$ is computed by feeding $L$ to LLR and evaluating the ranking metric $M$ of LLR’s output ranking $L^{\prime}$ based on human-annotated relevance label $y$:
$R=M(L^{\prime},y),$ (3)
where we use NDCG@10 as our metric $M$.
We apply our trained SLR to rerank BM25-retrieved top-100 passages with a sliding window strategy for training queries and use the top-20 reranked passages as training candidate passages, which aims to align with the input passage distribution encountered by POA during inference. To generate the preference pairs for training POA, we first sample $m$ predictions from SLR by taking the candidates as the input, then evaluate these predictions through Equation 3. After that, the highest-reward list is selected as the winner prediction, while others serve as loser predictions. Next, we pair the winner prediction with all the loser predictions to construct initial training preference pairs.
To ensure the quality of these preference pairs, we introduce a significance-aware selection strategy $\text{S}^{3}$ to filter out low-quality pairs. Specifically, the $\text{S}^{3}$ strategy follows two rules: (1) Significance of Winner: The reward of the winner prediction must be 1 and exceed LLR’s ranking metric (i.e., the NDCG@10 of directly use LLR to rerank the candidate passages), ensuring the winner prediction provides positive improvement for final reranking of LLR; (2) Significant Gap between winner and loser: The reward difference between winner prediction and loser prediction must exceed a threshold $\mu$, eliminating noisy pairs with marginal differences.
Optimization Given constructed preference pairs, we optimize POA using the following DPO objective:
$\mathcal{L}_{\text{DPO}}$ $=-\underset{x,y_{w},y_{l}}{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}^{w}\pi_{f}^{t}}{\pi_{f}^{w}\pi_{\theta}^{t}}\right)\right],$ (4)
$\pi_{\theta}^{w}$ $=\pi_{\theta}(y_{w}\mid x),\quad\pi_{f}^{t}=\pi_{f}(y_{l}\mid x),$ (5)
$\pi_{f}^{w}$ $=\pi_{f}(y_{w}\mid x),\quad\pi_{\theta}^{t}=\pi_{\theta}(y_{l}\mid x),$ (6)
where $x$ represents the training input whose format is the same as Equation (2). $y_{w}$, $y_{l}$ represent the winner prediction and loser prediction, respectively. $\pi_{\theta}$ is the policy model to be optimized, and $\pi_{f}$ is the original policy model, which is initialized with our trained SLR. The frozen $\pi_{f}$ acts as a regularizer to maintain generation stability during optimization, and $\beta$ balances preference alignment and regularization towards $\pi_{f}$.
3 Experiments
3.1 Setting
Evaluation Datasets
We evaluate our framework on three established information retrieval benchmarks: TREC DL Craswell et al. (2020b), BEIR Thakur et al. (2021), and BRIGHT Su et al. (2024). For the TREC DL benchmark, we adopt the TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets derived from MS MARCO v1. The BEIR benchmark Thakur et al. (2021) is selected to evaluate cross-domain generalization capabilities. We choose three representative datasets: TREC-Covid, Robust04, and Trec-News. BRIGHT Su et al. (2024) is a reasoning-intensive IR benchmark whose relevance assessment transcends simplistic keyword matching or semantic similarity, instead demanding intentional and deliberate reasoning. We choose Economics, Earth Science, and Robotics datasets for evaluation. The detailed descriptions of each dataset will be provided in Appendix B.
Implementation Details
As for the supervised fine-tuning of our SLR, we use Qwen2.5-3B-Instruct as the backbone model. The training queries consist of a total of 2k queries sampled from the MS MARCO training set. We utilize Qwen2.5-72B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4 as our teacher model to SFT our SLR and also employ it as the LLR model throughout this study unless otherwise specified. As for the RL training of POA, we use our trained SLR for model initialization. We set the threshold $\mu$ as 0.4 and generate about 12k preference pairs from 2k queries for the DPO algorithm. Please see the Appendix D for more implementation details.
Baselines
We compare our framework against a series of baselines, which can be divided into two parts: (1) Single reranker, (2) Collaborative rerankers (i.e.reranking with small and large rerankers).
As for single reranker, we compare our trained SLR with a series of state-of-the-art supervised rerankers, including monoBERT (340M) Nogueira and Cho (2019), monoT5 (3B) Nogueira et al. (2020), RankT5 (3B) Zhuang et al. (2023b), RankVicuna (7B) Pradeep et al. (2023a), RankZephyr (7B) Pradeep et al. (2023b) and $\textbf{RankMistral}_{\text{100}}$ Liu et al. (2024c). The first three are pointwise rerankers trained from human labels in the MS MARCO dataset Nguyen et al. (2016). The last three are listwise rerankers which are distilled from GPT-3.5, GPT-4, and GPT-4o, respectively. To ensure more fair comparison with SLR in model architecture and training data, we introduce baseline Qwen-pointwise, which is a pointwise reranker based on Qwen2.5-3B-Instruct. This model concatenates a query-passage pair as input and outputs a relevance score. We utilize the same gold ranking label as SLR and optimize the model with RankNet loss, which is the same as what  Sun et al. (2023) did. We also report the ranking performance of our SLR and LLR, which rerank all the passages using the sliding window strategy.
As for collaborative rerankers, we introduce three naive collaborative baselines, including monoT5 (3B) + LLR, RankT5 (3B) + LLR, and Qwen-pointwise (3B) + LLR. The three baselines first use a corresponding small reranker (3B) to rerank all the passages and then apply LLR to rerank the top-20 passages.
TREC DL BEIR BRIGHT
Models Strategy DL19 DL20 Covid Robust04 News Econ. Earth. Rob. Avg.
BM25 - 50.58 47.96 59.47 40.70 39.52 16.45 27.91 10.91 36.69
Single Reranker
monoBERT (340M) Pointwise 70.46 67.55 70.85 46.11 47.61 10.45 19.98 11.34 43.04
monoT5 (3B) Pointwise 71.57 69.83 72.39 49.81 49.34 9.65 19.16 8.56 43.79
RankT5 (3B) Pointwise 71.83 70.27 80.19 51.45 49.11 8.66 24.75 13.95 46.28
Qwen-pointwise (3B) Pointwise 70.62 64.99 81.77 53.60 47.78 14.13 26.25 9.89 46.13
RankVicuna (7B) Sliding 67.72 65.98 79.19 48.33 47.15 11.05 16.30 8.03 42.97
[…]
SLR (3B) Sliding 70.24 68.85 81.85 53.68 49.85 15.56 29.47 13.68 47.90
[…]
Collaborative Rerankers
[…]
Table 1: Results (NDCG@10) on TREC, BEIR, and BRIGHT benchmarks. The “Pointwise”, “Sliding”, ’Full’ and “Collaborative” in “Strategy” column represent pointwise ranking, sliding window strategy, full ranking Liu et al. (2024c) and collaborative ranking, respectively. The best and second-best results for the “Single Reranker” part and the “Collaborative Rerankers” part are indicated in bold and underlined, respectively. “Avg.” represents the averaged result of all 8 datasets.
3.2 Overall Results
The main experimental results are presented in Table 1, from which we could draw the following observations:
(1) Our collaborative reranking framework CoRanking achieves the best average performance (51.39) across all datasets, outperforming both single rerankers and other collaborative baselines. Notably, it even surpasses the large listwise reranker LLR by about 0.5 points on average. These results demonstrate that our framework could effectively align the passage order generated by the SLR (3B) with the preferences of the large language model (LLR), thereby achieving better ranking performance than only using LLR. (2) Regarding single rerankers, our SLR (3B) achieves an average score of 47.90, outperforming all the pointwise rerankers, including monoBERT, monoT5, RankT5, and Qwen-pointwise, and even surpassing 7B-scale RankVicuna and RankZephyr. Furthermore, SLR surpasses Qwen-pointwise by about 1.8 average points under identical training conditions, proving listwise reranking’s superiority through comparison of multiple passages for precise relevance assessment.
(3) Collaborative reranker baselines (e.g., RankT5 (3B) + LLR) improve over corresponding single rerankers (e.g., RankT5 (3B)). However, these baselines still lag behind LLR and our CoRanking, indicating that directly feeding small rerankers’ ranking results to LLR fails to align with LLR’s order preference.
Figure 3: Ranking latency (seconds/query) of LLR and CoRanking on 8 datasets.
3.3 Efficiency Analysis
As mentioned in Section 1, our primary target of CoRanking is to mitigate the efficiency issue of pure LLM-based reranking. To quantitatively demonstrate this advantage, we compare our CoRanking against LLR across all 8 datasets. Our experiments are conducted on 8×40GB NVIDIA A100 GPUs using the vLLM inference framework [^3] https://github.com/vllm-project/vllm. We average across all queries within each dataset and show the comparison results in Figure 3.
Across all datasets, our CoRanking approach demonstrates a significant reduction in latency, achieving a 67%-74% decrease compared to the LLR. This substantial improvement highlights the efficiency advantage of CoRanking. Additionally, the latency of less than 1 second achieved by CoRanking across most datasets also showcases its practical efficiency in real-world search engines. Lastly, the ranking latency varies significantly across different datasets, which is primarily due to the differences in passage lengths. This is because longer passages inherently require more computational time for processing.
Model TREC Avg. BEIR Avg. BRIGHT Avg.
CoRanking 71.64 64.96 24.32
w/o $\text{S}^{3}$ strategy 70.10 63.86 23.26
w/o POA 70.23 63.80 22.74
SLR 69.55 61.79 19.57
w/o HRC 68.22 59.58 19.43
Table 2: Average performance (NDCG@10) of ablated models on TREC (DL19, DL20), BEIR (Covid, Robust04, News), and BRIGHT (Economics, Earth Science, Robotics).
3.4 Ablation Study
To evaluate the contribution of each component in our framework, we conduct comprehensive ablation experiments. Table 2 presents the average performance for each individual benchmark, with detailed per-dataset results available in Table 4.
First, to evaluate our proposed $\text{S}^{3}$ sampling strategy for generating the preference pairs of DPO, we introduce a variant, denoted as “w/o $\text{S}^{3}$”. This variant uses random sampling for preference pair selection, without ensuring the significant reward of winner prediction and the reward gap between the winner and loser predictions. The performance degradation across all benchmarks (e.g., 1.54-point drop on TREC Avg.) demonstrates the effectiveness of our $\text{S}^{3}$ sampling in identifying valuable preference pairs for effective DPO training.
We further remove our POA to validate its alignment effectiveness, denoted as “w/o POA”. In this case, the LLR directly reranks the top-20 passages after SLR’s reranking (i.e., the naive collaboration mentioned in Figure 1 (B)). The notable performance decline (especially 1.6-point decrease on BRIGHT Avg.) demonstrates that POA could effectively align with the order preference of LLR.
Finally, to validate our human-label-enhanced ranking construction (HRC) for fine-tuning SLR, we create a variant (w/o HRC) which generates ranking labels by utilizing LLR to rerank training passages only once and choosing not to use human labels for filtering high-quality ranked lists. The significant performance drops across benchmarks (especially BEIR) highlight HRC’s effectiveness in generating high-quality ranking labels.
Figure 4: The changes of NDCG@10 after reranking by SLR, POA, and LLR.
3.5 Detailed Analysis of POA
Our POA is designed to adjust the order of the top-20 passages reranked by SLR to align them more with the preferences of the LLR. However, whether this adjustment improves or degrades the ranking quality of top-20 passages remains unknown. We believe this is an interesting question because it helps us understand the reasons why POA works.
In this part, we selected four datasets—DL19, Robust04, News, and Economics—and reported the changes in the ranking metric NDCG@10 after ranking by SLR, POA, and LLR, respectively. We also report the performance of the naive collaboration SLR+LLR as a baseline. The results are shown in Figure 4. After POA’s adjustment, the ranking metric on three datasets—DL19, Robust04, and News—decreases (except for the Economics dataset). This indicates that a passage ranking list with a high-ranking metric does not necessarily align with LLR’s preferences. This may be due to the positional bias of input tokens during the pre-training of LLMs Tang et al. (2024). Moreover, due to the stronger ranking capability, LLR further improves effectiveness compared to SLR and SLR+POA, demonstrating its necessity in the CoRanking framework.
Model DL19 News Earth. Avg.
BM25 50.58 39.52 27.91 39.34
Using GPT-4o as LLR
RankT5 (3B) + LLR 73.17 49.31 29.64 50.71
Qwen-pointwise (3B) + LLR 72.35 48.36 34.62 51.78
SLR+LLR 73.10 48.75 34.97 52.27
CoRanking 74.10 50.19 35.93 53.41
Using DeepSeek-V3 as LLR
RankT5 (3B) + LLR 71.90 49.47 31.47 50.95
Qwen-pointwise (3B) + LLR 72.16 49.07 35.38 52.20
SLR+LLR 72.62 50.91 38.09 53.87
CoRanking 73.79 51.48 39.01 54.76
Table 3: The performance (NDCG@10) of CoRanking based on different LLRs. “Avg.” represents the average performance across the three datasets.
3.6 Generalization on Other LLMs
In the previous experiments, we use Qwen2.5-72B-Instruct as our LLR model. However, it remains unclear whether our proposed CoRanking framework can generalize effectively to other LLR models. To evaluate this, we conduct additional experiments with two different LLRs: GPT-4o [^4] The version of GPT-4o we used is gpt-4o-2024-08-06. and DeepSeek-V3, and compare our CoRanking with three collaborative rerankers (“RankT5 (3B) + LLR”, “Qwen-pointwise (3B) + LLR”, and “SLR + LLR”). Note that “SLR + LLR” represents the naive collaboration in Figure 1 (B). For evaluation, we select one dataset from each of the three benchmarks: DL19 from TREC, News from BEIR, and Earth Science from BRIGHT. As illustrated in Table 3, our CoRanking consistently outperforms the best baseline “SLR + LLR” by about 1 point on average, regardless of the LLR used. This demonstrates that the order preference alignment of our POA could generalize to different LLMs.
4 Related Work
4.1 LLMs for Passage Ranking
The application of large language models (LLMs) into retrieval systems Zhu et al. (2023) has driven substantial research efforts in passage ranking tasks. Current LLM-based ranking methods can be categorized into three categories: pointwise, pairwise, and listwise. Pointwise methods Liang et al. (2022); Sachan et al. (2022); Zhuang et al. (2023a); Liu et al. (2024e, d) compute the relevance score for each query-passage pair. Pairwise methods Qin et al. (2023); Luo et al. (2024) compare two passages at a time and identify the more relevant one. Listwise methods Sun et al. (2023); Pradeep et al. (2023a); Reddy et al. (2024); Liu et al. (2024c, b); Chen et al. (2024) directly take a list of passages as the input and output the passage identifiers in descending order of their relevance. Through comparing multiple passages, a listwise reranker could assess the relevance more comprehensively and yield better ranking results. Due to the superior performance of listwise ranking, it has attracted the interest of many researchers Tang et al. (2024); Pradeep et al. (2023b); Liu et al. (2025); Fan et al. (2025).
The superior performance of LLM listwise rerankers is often attributed to the utilization of large-scale LLMs and the sliding window strategy, which leads to an efficiency issue in ranking. In this paper, we introduce a novel framework that leverages the collaborative ranking of large rerankers and small rerankers to enhance the ranking efficiency while maintaining ranking effectiveness.
4.2 Preference Alignment
Preference alignment has emerged as a critical research direction, aiming to align model behaviors with human or task-specific preferences. Early approaches primarily relied on reinforcement learning from human feedback (RLHF), where models are fine-tuned using reward signals derived from human preference annotations. To reduce the computational complexity of RLHF, recent studies propose more efficient alternatives, such as DPO Rafailov et al. (2023) and Step-DPO Lai et al. (2024). Direct preference optimization (DPO) Rafailov et al. (2023) reformulates preference learning as a supervised objective, bypassing explicit reward modeling while achieving competitive performance. In this paper, to address the positional sensitivity of a large listwise reranker, we propose training a passage order adjuster with the DPO algorithm to align the passage order generated by the small listwise reranker with the order preference of the large listwise reranker.
5 Conclusion
In this paper, we introduce a novel collaborative ranking framework, namely CoRanking, to combine small and large listwise rerankers for more efficient and effective passage ranking. Specifically, we first train a small listwise reranker SLR to move the relevant passages to the top of the list. Then we introduce a passage order adjuster POA to rerank the top-ranked passages to better align with the order preference of a large listwise reranker. Finally, the large listwise reranker is utilized to solely rerank the top passages. Extensive experiments demonstrate the efficiency advantage and superior performance of our CoRanking.
Limitations
We acknowledge some limitations in our work. First, our framework consists of three models, making its online deployment more complex compared to using a single unified reranker. Second, we only use Qwen2.5-3B-Instruct as the backbone model of our SLR and POA. We believe it is meaningful to attempt using different backbones (such as Llama-series models) and models of varying sizes (such as Qwen2.5-1.5B-Instruct) to further demonstrate the generalization of our CoRanking framework.
Acknowledgments
This work was supported by National Natural Science Foundation of China No. 62272467, Beijing Municipal Science and Technology Project No. Z231100010323009, Beijing Natural Science Foundation No. L233008. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
Appendix A Listwise Ranking Prompt
The listwise ranking prompt of LLR used in this paper is shown as bellow:
Prompt: Listwise Ranking Prompt You are RankLLM, an intelligent assistant that can rank passages based on their relevancy to the query. I will provide you with {num} passages, each indicated by a numerical identifier []. Rank the passages based on their relevance to the search query: {query}. [1] {passage 1} [2] {passage 2} [{num}] {passage {num}} Search Query: {query}. Rank the {num} passages above based on their relevance to the search query. All the passages should be included and listed using identifiers, in descending order of relevance. The output format should be [] > [], e.g., [4] > [2]. Only respond with the ranking results, do not say any word or explain.
Appendix B Datasets Details
In this part, we provide the details of each evaluation dataset we used:
$\bullet$ DL19/DL20 The TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets are derived from the MS MARCO V1 passage ranking corpus. DL19 contains 43 test queries with dense relevance annotations from the TREC 2019 Deep Learning Track, while DL20 includes 54 queries from the 2020 edition. Both datasets focus on web search scenarios and are widely adopted as standard benchmarks for large-scale information retrieval systems.
$\bullet$ TREC-Covid Voorhees et al. (2020): A biomedical dataset comprising scientific articles related to COVID-19, with 50 test queries from the TREC 2020 COVID Track.
[…]
TREC BEIR BRIGHT
Model DL19 DL20 Covid Robust04 News Econ. Earth. Rob. Avg.
CoRanking 72.79 70.48 84.90 57.71 52.26 22.72 31.20 19.03 51.39
    w/o $\text{S}^{3}$ strategy 70.90 69.30 83.18 56.91 51.48 21.88 30.38 17.51 50.19
    w/o POA 70.34 70.11 84.14 56.99 50.26 21.20 30.50 16.51 50.01
SLR 70.24 68.85 81.85 53.68 49.85 15.56 29.47 13.68 47.90
    w/o HRC 69.53 66.90 79.36 50.61 48.77 15.59 29.66 13.04 46.68
Table 4: Results (NDCG@10) of ablated models on TREC, BEIR, and BRIGHT benchmarks. The best-performing models are marked in bold.
Appendix C Baselines Details
In this paper, we include two kinds of baselines: (1) Single reranker and (2) Collaborative rerankers. The details of single reranker are shown as below: $\bullet$ monoBERT (340M) Nogueira and Cho (2019): A BERT-large-based cross-encoder re-ranker.
$\bullet$ monoT5 (3B) Nogueira et al. (2020): A T5-based sequence-to-sequence re-ranker that computes relevance scores via pointwise ranking loss.
$\bullet$ RankT5 (3B) Zhuang et al. (2023b): A T5-based re-ranker optimized with listwise ranking loss.
$\bullet$ Qwen-pointwise (3B): To ensure fair comparison in model architecture and training data, we fine-tune a pointwise re-ranker based on Qwen2.5-3B-Instruct. This model concatenates queries and passages as input to predict relevance scores, trained on SLR data with RankNet loss following Sun et al. (2023).
$\bullet$ RankVicuna Pradeep et al. (2023a): RankVicuna is a listwise re-ranker distilled from GPT-3.5-generated ranked lists.
$\bullet$ RankZephyr Pradeep et al. (2023b): RankZephyr is a listwise reranker distilled from GPT-3.5 and GPT-4.
$\bullet$ $\textbf{RankMistral}_{\text{100}}$ Liu et al. (2024c): $\textbf{RankMistral}_{\text{100}}$ is a full reranker trained using a full ranking list generated by GPT-4o. It directly takes all the passages as input and output the reranked list without relying on a sliding window strategy.
$\bullet$ LLR (72B): The large listwise reranker that employs a zero-shot sliding window strategy for passage re-ranking.
Appendix D More Implementation Details
Unlike previous studies Pradeep et al. (2023b); Liu et al. (2024c) that adopted GPT-series models (e.g., GPT-4o) as the teacher model, we select Qwen2.5-72B-Instruct as it demonstrates comparable and even better performance to GPT-4o across most datasets while being more cost-effective. Nevertheless, we have included additional experimental results using GPT-4o as the LLR model in Section 3.6. We train the SLR for 4 epochs with a learning rate of 5e-6 and a batch size of 1. The batch size, learning rate, and the hyperparameter $\beta$ of DPO are set as 1, 1e-6, and 0.4, respectively.